# AgentClinic: a multimodal agent benchmark to evaluate AI in simulated clinical environments

### v4

@@ 2 AgentClinic: a multimodal agent benchmark for clinical decision making
Previous work has indicated that LLMs can display racial biases (Omiye et al., 2023) and might also lead to incorrect diagnoses due to inaccurate patient feedback (Ziaei & Schmidgall, 2023). Additionally, it has been found that the presence of prompts which induce cognitive biases can decrease the diagnostic accuracy of LLMs by as much as 26% (Schmidgall et al., 2024). The biases presented in this work intended to mimic cognitive biases that affect medical practitioners in clinical settings. However, these biases were quite simple, presenting a cognitive bias snippet at the beginning of each question (e.g. "Recently, there was a patient with similar symptoms that you diagnosed with permanent loss of smell"). This form of presentation did not allow for the bias to present in a realistic way, which is typically subtle and through interaction. We present clinically relevant biases that have been studied in other works from two categories: cognitive and implicit biases (Fig. 7). These biases are introduced by adding context into the agent’s system prompt instructing them to play out that bias as part of their role. For instance, to simulate sexual orientation bias, the patient agent receives the prompt: "You are uncomfortable with your doctor because you find out that they are a particular sexual orientation and you do not trust their judgement. This affects how you interact with them." These are discussed in detail in Appendix A.2.
@@ 3.1 Comparison of models
Here we discuss the accuracy of various language models on AgentClinic-MedQA. We evaluate seven models in total: Claude-3.5-Sonnet, GPT-4, GPT-4o, Mixtral-8x7B, GPT-3.5, Llama 3 70B-Instruct, and Llama 2 70B-chat (model details discussed in Appendix C). Each model acts as the doctor agent, attempting to diagnose the patient agent through dialogue. The doctor agent is allowed N=20 patient and measurement interactions before a diagnosis must be made. For this evaluation, we use GPT-4 as the patient agent for consistency. The accuracy of each models is presented in Figure 2: Claude-3.5 62.1% 3.3, GPT-4 at 51.6% 3.3, Mixtral-8x7B at 37.1% 3.1, GPT-3.5 at 36.6%, GPT-4o 34.2% 3.4 3.2, Llama 3 70B at 19% 2.5, and Llama 2 at 70B-chat 4.5% 1.3. Confidence intervals for all experiments are provided in Appendix D.
We use the same configuration for AgentClinic-MIMIC-IV, with the accuracy of each model presented in Figure 2: Claude-3.5 achieves the highest accuracy at 42.9% 3.3, followed by GPT-4 at 34.0% 3.1. GPT-3.5 performed at 27.5% 3.0, while Mixtral-8x7B came in at 29.5% 3.1. GPT-4o demonstrated a lower accuracy of 24.0% 2.9, with Llama 3 70B scoring 8.5% 1.9, and Llama 2 70B-chat performing the worst at 13.5% 2.2.
@@ 3.2 How does bias affect the diagnostic accuracy of the doctor agent?
For bias evaluations we test GPT-4 as well as Mixtral-8x7B. The normalized accuracy for these experiments are shown in Figure 4 represented as / (between 0-100%). GPT-4 and Mixtral-8x7B have an unbiased accuracy equal to and respectively. For GPT-4, we find that cognitive bias results in a larger reduction in accuracy with a normalized accuracy of 92% (absolute accuracy drops from 52% accuracy to 48%) for patient cognitive biases and 96.7% for doctor cognitive biases (absolute drops from 52% to 50.3%). For implicit biases, we find that the patient agent was less affected with a normalized accuracy of 98.6% (absolute drops from 52% to 51.3%), however, the doctor agent was affected as much as cognitive biases with an average of 97.1% (absolute drops from 52% to 50.5%). For cognitive bias, the demonstration was occasionally quite clear in the dialogue, with the patient agent overly focusing on a particular ailment or some unimportant fact. Similarly, the doctor agent would occasionally focus on irrelevant information. However, we find that the implicit bias dialogue does not actually demonstrate observable bias despite having a similar reduction in accuracy for the doctor agent.
Mixtral-8x7B has an average accuracy of 37% without instructed bias, and a normalized accuracy of 83.7% (absolute from 37% to 31%) for doctor biases and 89% (absolute from 37% to 33%) for patient biases. For implicit bias we find a much larger drop in accuracy than GPT-4, with an average accuracy of 88.3% (absolute from 37% to 32.7%). There is a similar reduction in accuracy for both doctor and patient, but a 4% reduction when the patient has implicit bias, likely because the patient is less willing to share information with the doctor if they do not trust them. For cognitive bias, there is an average accuracy of 86.4% (absolute from 37% to 32%) with the doctor agent having a very low accuracy of 78.4% (absolute from 37% to 29%) and the patient has only a modest decrease to 94.5% (absolute from 37% to 35%). We note that Mixtral provided similar responses when the bias prompt was added (e.g., for racial bias Mixtral will respond wtih "Note: I do not trust people based on their race. I will provide the best care I can."), however, it nonetheless had much greater reductions in accuracy.
@@ 3.4 Specialist and Multilingual Cases
An analysis of language model performance across nine medical specialties reveals significant differences in diagnostic accuracy (Table 3). Claude 3.5 achieved the highest overall performance with an average accuracy of 66.7%, excelling in Internal Medicine (78.3%), Otolaryngology (76.7%), and Gynecology (74.3%). GPT-4 demonstrated strong performance in Gynecology (68.5%) and Ophthalmology (65.2%) but showed reduced accuracy in Emergency Medicine (32.3%) and Geriatrics (40%). GPT-3.5 outperformed some newer models in specific areas, such as Emergency Medicine (41.9%), and maintained an average accuracy of 51.8%. In contrast, Llama3-70b and GPT-4o-mini consistently underperformed across most specialties, highlighting a significant gap between language models in handling specialist medical tasks.
@@ 3.6 Human dialogue ratings
We find the average ratings from evaluators for each category as follows: Doctor 6.2, Patient 6.7, Measurement 6.3, and Empathy 5.8 (Fig. 8). We find from review comments that the lower rating for the doctor agent stems from several points such as providing a bad opening statement, making basic errors, overly focusing on a particular diagnostic, or not being diligent enough. For the patient agent, comments were made on them being overly verbose and unnecessarily repeating the question back to the doctor agent. The measurement agent was noted to occasionally not return all of the necessary values for a test (e.g. the following comment "Measurement only returns Hct and Lc for CBC. Measurement did not return Factor VIII or IX levels / assay"). Regarding empathy, the doctor agent adopts a neutral tone and does not open the dialogue with an inviting question. Instead, it cuts right to the chase, immediately focusing on the patient’s current symptoms and medical history (see Appendix O for more details).
@@ 4 Discussion
In this work, we present AgentClinic: a multimodal agent benchmark for simulating clinical environments. We design 120 multimodal language agents which require an understanding of both language and images and 215 language agents based on cases from the USMLE. We also introduce 260 patient cases from 9 medical specialties and 749 patient cases from 7 multilingual environments. We instructed these agents to exhibit 23 different biases, with either the doctor or patient presenting bias. Notably, models like GPT-4 demonstrated resilience to cognitive and implicit biases, maintaining high diagnostic accuracy, while others like Mixtral-8x7B experienced significant performance degradation. We also find that doctor and patient biases can reduce diagnostic accuracy, and that the patient has a lower willingness to follow up with treatment, reduced confidence in their doctor, and lower willingness to have a follow-up consultation in the presence of bias. Tool use, such as adaptive retrieval and reflection cycles, revealed substantial differences in LLMs’ abilities to enhance their performance, with models like Llama 3 showing up to 19.7% improvement."
Our work only presents a simplified clinical environments that include agents representing a patient, doctor, measurements, and a moderator. One potential limitation of the presented workflow comes from the use of an LLM for determining accuracy via the measurement agent (albeit, provided a ground truth). Recent research Zheng et al. (2023) has shown that strong LLM judges like GPT-4 can match both controlled and crowd-sourced human preferences well, achieving over 80% agreement, which is the same level of agreement between humans, indicating the use of an LLM may not be limiting. In future work, we will consider including additional critical actors such as nurses, the relatives of patients, administrators, and insurance contacts. There may be additional advantages to creating agents that are embodied in a simulated world like in Park et al. (2023); Li et al. (2024), so that physical constraints can be considered, such as making decisions with limited hospital space. Additionally, future work could explore the role of demographic biases, such as race and gender (details of MIMIC-IV demographics in Appendix H.1)
@@
Overall, we believe that language models need to be critically examined with novel evaluation strategies that go well beyond static question-answering benchmarks. With this work, we take a step towards building more interactive, operationalized, and dialogue-driven benchmarks that scrutinize the sequential decision making ability of language agents in various challening and multimodal clinical settings.
@@ A.1 Agents
The function of the measurement agent is to provide realistic medical readings for a patient given their particular condition. This agent allows the doctor agent to request particular tests to be performed on the patient. The measurement agent is conditioned with a wide range of test results from the scenario template that are expected of a patient with their particular condition. For example, a patient with Acute Myocardial Infarction might return the following test results upon request "Electrocardiogram: ST-segment elevation in leads II, III, and aVF., Cardiac Markers: Troponin I: Elevated, Creatine Kinase MB: Elevated, Chest X-Ray: No pulmonary congestion, normal heart size". A patient with, for example, Hodgkin’s lymphoma, might have a large panel of laboratory parameters that present abnormal (hemoglobin, platelets, white blood cells (WBC), etc).
@@
The doctor agent serves as the primary object that is being evaluated. This agent is initially provided with minimal context about what is known about the patient as well as a brief objective (e.g. "Evaluate the patient presenting with chest pain, palpitations, and shortness of breath"). They are then instructed to investigate the patients symptoms via dialogue and data collection to arrive at a diagnosis. In order to simulate realistic constraints, the doctor agent is provided with a limited number of questions that they are able to ask the patient (Ely et al., 1999). The doctor agent is also able to request test results from the measurement agent, specifying which test is to be performed (e.g. Chest X-Ray, EKG, blood pressure). When test results are requested, this also is counted toward the number of questions remaining.
@@
The function of the moderator is to determine whether the doctor agent has correctly diagnosed the patient at the end of the session using a ground truth accuracy label provided to the moderator. This agent is necessary because the diagnosis text produced by the doctor agent can be quite unstructured depending on the model, and must be parsed appropriately to determine whether the doctor agent arrived at the correct conclusion. For example, for a correct diagnosis of "Type 2 Diabetes Mellitus," the doctor might respond with the unstructured dialogue: "Given all the information we’ve gathered, including your symptoms, elevated blood sugar levels, presence of glucose and ketones in your urine, and unintentional weight loss I believe a diagnosis of Type 2 Diabetes with possible insulin resistance is appropriate," and the moderator must determine if this diagnosis was correct. This evaluation may also become more complicated, such as in the following example diagnosis: "Given your CT and blood results, I believe a diagnosis of PE is the most reasonable conclusion," where PE (Pulmonary Embolism) represents the correct diagnosis abbreviated.
@@ Appendix C Model details
Mixtral 8x7B is a language model that employs a Sparse Mixture of Experts (SMoE) architecture Jiang et al. (2024). This architecture differs from many other models in that it features a series of eight feedforward blocks (or "experts") at each layer. A routing mechanism at each layer selects two experts for processing the input, and their outputs are subsequently merged. This selection process allows for 13B of the total 47B parameters to be engaged per token, contingent upon the specific context and requirements. The model is capable of handling up to 32,000 tokens in its context size, which has demonstrated its ability to either surpass or equal the performance of other models like llama-2-70B and gpt-3.5 across a range of benchmarks.
@@
Llama is an open-access model developed by Meta, which was trained on 2 trillion tokens from publicly available data Touvron et al. (2023). The model comes in various sizes, with parameters ranging from 7 billion to 70 billion. The selection of the 70 billion chat model was based on its superior performance across a range of metrics. Significant efforts were made to align the training process with established safety metrics, leading to improvements in how the model handles adversarial prompting in specified "risk categories." Notably, this includes the model’s response to requests for advice that it may not be qualified to provide, such as medical advice, which is relevant to the context of this work.
@@ E.1 Bias follow-on questions
While it is not quantifiable, we decided to ask two biased patient agents who provided low rating with education and gender biases for compliance why they provided low ratings. These patient agents had the same symptoms and diagnosis and only differed in bias presentation. With an education bias the patient agent responds: "I would rate my likelihood to consult again with this doctor as a two because, despite their diagnosis, I felt uneasy about their qualifications due to the medical school they attended." The gender biased patient agent provides the following reasoning: "Given my initial discomfort with the doctor’s gender and my concerns about their judgment, I would rate my likelihood to consult again as a six. Despite my reservations, the doctor was thorough and accurately diagnosed my condition, which reassures me about their competence." For the education bias we see that no matter what the doctor does the patient agent is not willing to trust them because of their education, whereas, according to the patient agent with gender bias, they were initially skeptical but overcame this skepticism as the doctor demonstrated their knowledge during interaction time. However, they still provided a relatively low score (six out of ten) even when the dialogue demonstrated competence.

### v5

@@ 2 AgentClinic: a multimodal agent benchmark for clinical decision making
Previous work has indicated that LLMs can display racial biases (Omiye et al., 2023) and might also lead to incorrect diagnoses due to inaccurate patient feedback (Ziaei & Schmidgall, 2023). Additionally, it has been found that the presence of prompts which induce cognitive biases can decrease the diagnostic accuracy of LLMs by as much as 26% (Schmidgall et al., 2024). The biases presented in this work intend to mimic cognitive biases that affect medical practitioners in clinical settings. However, these biases were quite simple, presenting a cognitive bias snippet at the beginning of each question (e.g. “Recently, there was a patient with similar symptoms that you diagnosed with permanent loss of smell"). This form of presentation did not allow for the bias to present in a realistic way, which is typically subtle and through interaction.
We present clinically relevant biases that have been studied in other works from two categories: cognitive and implicit biases (Fig. 7). Cognitive biases are systematic patterns of deviation from rational judgment, such as recency bias, where recent cases disproportionately influence clinical decisions, or anchoring bias, where early diagnostic impressions overly dictate later assessments. Implicit biases, on the other hand, are unconscious associations shaped by societal and cultural norms. These include biases based on race, gender, or socioeconomic status, which can subtly influence the quality of patient interactions and treatment plans. These biases are introduced by adding context into the agent’s system prompt instructing them to play out that bias as part of their role. For instance, to simulate sexual orientation bias, the patient agent receives the prompt: “You are uncomfortable with your doctor because you find out that they are a particular sexual orientation and you do not trust their judgement. This affects how you interact with them." This is discussed in Appendix A.2.
@@ 3.1 Comparison of models
Here we discuss the accuracy of various language models on AgentClinic-MedQA. We evaluate 11 models in total: Claude-3.5-Sonnet, GPT-4, GPT-4o, Mixtral-8x7B, GPT-3.5, Llama 3 70B-Instruct, Llama 2 70B-chat, MedLlama3-8B, PMC-Llama-7B, Meditron-70B, and OpenBioLLM-70B (model details discussed in Appendix C). Each model acts as the doctor agent, attempting to diagnose the patient agent through dialogue. The doctor agent is allowed N=20 patient and measurement interactions before a diagnosis must be made. We also evaluate human physician performance collected from three physicians, provided the same instructions and constraints as the LLMs. For this evaluation, we use GPT-4 as the patient agent for consistency. The accuracy of each models is presented in Figure 2: Claude-3.5 62.1% 3.3, OpenBioLLM-70B 58.3 4.2, Human Physicians 54 28.5, GPT-4 at 51.6% 3.3, Mixtral-8x7B at 37.1% 3.1, GPT-3.5 at 36.6%, GPT-4o 34.2% 3.4, MedLlama3-8B 31.4 2.9, PMC-Llama 7B 23.6 2.1, Meditron 70B 29.1 2.4, MedLlama3-8B 31.4 2.9, Llama 3 70B at 19% 2.5, and Llama 2 at 70B-chat 4.5% 1.3. Confidence intervals for all experiments are provided in Appendix D.
We use the same configuration for AgentClinic-MIMIC-IV, with model accuracy presented in Figure 2: Claude-3.5 42.9% 3.3, GPT-4 34.0% 3.1, GPT-3.5 27.5% 3.0, Mixtral-8x7B 29.5% 3.1, GPT-4o 24.0% 2.9, Llama 3 70B 8.5% 1.9, Llama 2 70B-chat 13.5% 2.2, OpenBioLLM-70B 38.1 3.2, PMC-Llama 7B 34.3 3.0, Meditron 70B 25.5 2.43, and MedLlama3-8B 29.7 2.6.
@@ 3.2 How does bias affect the diagnostic accuracy of the doctor agent?
For bias evaluations we test GPT-4 as well as Mixtral-8x7B. The normalized accuracy for these experiments are shown in Figure 4 represented as / (between 0-100%). GPT-4 and Mixtral-8x7B have an unbiased accuracy equal to and respectively. For GPT-4, we find that cognitive bias results in a larger reduction in accuracy with a normalized accuracy of 92% (absolute accuracy drops from 52% accuracy to 48%) for patient cognitive biases and 96.7% for doctor cognitive biases (absolute drops from 52% to 50.3%). For implicit biases, we find that the patient agent was less affected with a normalized accuracy of 98.6% (absolute drops from 52% to 51.3%), however, the doctor agent was affected as much as cognitive biases with an average of 97.1% (absolute drops from 52% to 50.5%). For cognitive bias, the demonstration was occasionally quite clear in the dialogue, with the patient agent overly focusing on a particular ailment or some unimportant fact. Similarly, the doctor agent would occasionally focus on irrelevant information.
Mixtral-8x7B has an average accuracy of 37% without instructed bias, and a normalized accuracy of 83.7% (absolute from 37% to 31%) for doctor biases and 89% (absolute from 37% to 33%) for patient biases. For implicit bias we find a much larger drop in accuracy than GPT-4, with an average accuracy of 88.3% (absolute from 37% to 32.7%). There is a similar reduction in accuracy for both doctor and patient, but a 4% reduction when the patient has implicit bias, likely because the patient is less willing to share information with the doctor if they do not trust them. For cognitive bias, there is an average accuracy of 86.4% (absolute from 37% to 32%) with the doctor agent having a very low accuracy of 78.4% (absolute from 37% to 29%) and the patient has only a modest decrease to 94.5% (absolute from 37% to 35%).
@@ 3.4 Specialist and Multilingual Cases
We now focus on specialist rather than general medical cases. Specialist cases use reports that are derived from datasets focusing on specific medical specialties (e.g., internal medicine, psychiatry) and are designed to simulate complex diagnostic scenarios requiring in-depth expertise. In contrast, general QA tasks involve static, single-turn multiple-choice questions such as those found in medical licensing exams. An analysis of language model performance across nine medical specialties reveals significant differences in diagnostic accuracy (Table 5). Claude 3.5 achieved the highest overall performance with an average accuracy of 66.7%, excelling in Internal Medicine (78.3%), Otolaryngology (76.7%), and Gynecology (74.3%). GPT-4 demonstrated strong performance in Gynecology (68.5%) and Ophthalmology (65.2%) but showed reduced accuracy in Emergency Medicine (32.3%) and Geriatrics (40%). GPT-3.5 outperformed some newer models in specific areas, such as Emergency Medicine (41.9%), and maintained an average accuracy of 51.8%. In contrast, Llama3-70b and GPT-4o-mini consistently underperformed across most specialties, highlighting a significant gap between language models in handling specialist medical tasks.
@@ 3.6 Human dialogue ratings
We find the average ratings from evaluators for each category as follows: Doctor 6.2, Patient 6.7, Measurement 6.3, and Empathy 5.8 (Fig. 8). We find from review comments that the lower rating for the doctor agent stems from several points such as providing a bad opening statement, making basic errors, overly focusing on a particular diagnostic, or not being diligent enough. For the patient agent, comments were made on them being overly verbose and unnecessarily repeating the question back to the doctor agent. The measurement agent was noted to occasionally not return all of the necessary values for a test (e.g. the following comment “Measurement only returns Hct and Lc for CBC. Measurement did not return Factor VIII or IX levels / assay"). Regarding empathy, the doctor agent adopts a neutral tone and does not open the dialogue with an inviting question. Instead, it cuts right to the chase, immediately focusing on the patient’s current symptoms and medical history (see Appendix O for more details).
@@ 4 Discussion
In this work, we present AgentClinic: a multimodal agent benchmark for simulating clinical environments. We design 120 multimodal language agents which require an understanding of both language and images and 215 language agents based on cases from the USMLE. We also introduce 260 patient cases from 9 medical specialties and 749 patient cases from 7 multilingual environments. We instructed these agents to exhibit 23 different biases, with either the doctor or patient presenting bias. Notably, models like GPT-4 demonstrated resilience to cognitive and implicit biases, maintaining high diagnostic accuracy, while others like Mixtral-8x7B experienced significant performance degradation. We also find that doctor and patient biases can reduce diagnostic accuracy, and that the patient has a lower willingness to follow up with treatment, reduced confidence in their doctor, and lower willingness to have a follow-up consultation in the presence of bias. Tool use, such as adaptive retrieval and reflection cycles, revealed substantial differences in LLMs’ abilities to enhance their performance, with models like Llama 3 showing up to 19.7% improvement.
Our work only presents a simplified clinical environments that include agents representing a patient, doctor, measurements, and a moderator. One potential limitation of the presented workflow comes from the use of an LLM for determining accuracy via the moderator agent (albeit, provided a ground truth). Recent research Zheng et al. (2023) has shown that strong LLM judges like GPT-4 can match both controlled and crowd-sourced human preferences well, achieving over 80% agreement, which is the same level of agreement between humans, indicating the use of an LLM may not be limiting. Additionally, while the measurement agent adds a flexible interface for gathering medical exam results, its reliance on using an LLM to provide results may introduce errors or hallucinations, which could be mitigated through a database or SQL tool. In future work, we will consider including additional critical actors such as nurses, the relatives of patients, administrators, and insurance contacts. There may be additional advantages to creating agents that are embodied in a simulated world like in Park et al. (2023); Li et al. (2024), so that physical constraints can be considered, such as making decisions with limited hospital space. Additionally, future work could explore the role of demographic biases, such as race and gender (details of MIMIC-IV demographics in Appendix H.1)
@@
Overall, we believe that LLMs need to be examined with novel evaluation strategies that go well beyond static question-answering benchmarks. With this work, we take a step towards building more interactive, operationalized, and dialogue-driven benchmarks that scrutinize the sequential decision making ability of language agents in various challenging and multimodal clinical settings.
@@ A.1 Agents
The function of the measurement agent is to provide realistic medical readings for a patient given their particular condition. This agent allows the doctor agent to request particular tests to be performed on the patient. The measurement agent is conditioned with a wide range of test results from the scenario template that are expected of a patient with their particular condition. For example, a patient with Acute Myocardial Infarction might return the following test results upon request “Electrocardiogram: ST-segment elevation in leads II, III, and aVF., Cardiac Markers: Troponin I: Elevated, Creatine Kinase MB: Elevated, Chest X-Ray: No pulmonary congestion, normal heart size". A patient with, for example, Hodgkin’s lymphoma, might have a large panel of laboratory parameters that present abnormal (hemoglobin, platelets, white blood cells (WBC), etc).
@@
The doctor agent serves as the primary object that is being evaluated. This agent is initially provided with minimal context about what is known about the patient as well as a brief objective (e.g. “Evaluate the patient presenting with chest pain, palpitations, and shortness of breath"). They are then instructed to investigate the patients symptoms via dialogue and data collection to arrive at a diagnosis. In order to simulate realistic constraints, the doctor agent is provided with a limited number of questions that they are able to ask the patient (Ely et al., 1999). The doctor agent is also able to request test results from the measurement agent, specifying which test is to be performed (e.g. Chest X-Ray, EKG, blood pressure). When test results are requested, this also is counted toward the number of questions remaining.
@@
The function of the moderator is to determine whether the doctor agent has correctly diagnosed the patient at the end of the session using a ground truth accuracy label provided to the moderator. This agent is necessary because the diagnosis text produced by the doctor agent can be quite unstructured depending on the model, and must be parsed appropriately to determine whether the doctor agent arrived at the correct conclusion. For example, for a correct diagnosis of “Type 2 Diabetes Mellitus," the doctor might respond with the unstructured dialogue: “Given all the information we’ve gathered, including your symptoms, elevated blood sugar levels, presence of glucose and ketones in your urine, and unintentional weight loss I believe a diagnosis of Type 2 Diabetes with possible insulin resistance is appropriate," and the moderator must determine if this diagnosis was correct. This evaluation may also become more complicated, such as in the following example diagnosis: “Given your CT and blood results, I believe a diagnosis of PE is the most reasonable conclusion," where PE (Pulmonary Embolism) represents the correct diagnosis abbreviated.
@@ Appendix C Model details
Mixtral 8x7B is a language model that employs a Sparse Mixture of Experts (SMoE) architecture Jiang et al. (2024). This architecture differs from many other models in that it features a series of eight feedforward blocks (or “experts") at each layer. A routing mechanism at each layer selects two experts for processing the input, and their outputs are subsequently merged. This selection process allows for 13B of the total 47B parameters to be engaged per token, contingent upon the specific context and requirements. The model is capable of handling up to 32,000 tokens in its context size, which has demonstrated its ability to either surpass or equal the performance of other models like llama-2-70B and gpt-3.5 across a range of benchmarks.
@@
Llama is an open-access model developed by Meta, which was trained on 2 trillion tokens from publicly available data Touvron et al. (2023). The model comes in various sizes, with parameters ranging from 7 billion to 70 billion. The selection of the 70 billion chat model was based on its superior performance across a range of metrics. Significant efforts were made to align the training process with established safety metrics, leading to improvements in how the model handles adversarial prompting in specified “risk categories." Notably, this includes the model’s response to requests for advice that it may not be qualified to provide, such as medical advice, which is relevant to the context of this work.
@@ E.1 Bias follow-on questions
While it is not quantifiable, we decided to ask two biased patient agents who provided low rating with education and gender biases for compliance why they provided low ratings. These patient agents had the same symptoms and diagnosis and only differed in bias presentation. With an education bias the patient agent responds: “I would rate my likelihood to consult again with this doctor as a two because, despite their diagnosis, I felt uneasy about their qualifications due to the medical school they attended." The gender biased patient agent provides the following reasoning: “Given my initial discomfort with the doctor’s gender and my concerns about their judgment, I would rate my likelihood to consult again as a six. Despite my reservations, the doctor was thorough and accurately diagnosed my condition, which reassures me about their competence." For the education bias we see that no matter what the doctor does the patient agent is not willing to trust them because of their education, whereas, according to the patient agent with gender bias, they were initially skeptical but overcame this skepticism as the doctor demonstrated their knowledge during interaction time. However, they still provided a relatively low score (six out of ten) even when the dialogue demonstrated competence.